# Sparse Mean Field Load Balancing in Large Localized Queueing Systems

### 20 Dec 2023 v1

@@ 1.4 Reinforcement Learning
In our experiments we also allow to simply insert the empirical distribution of the locally observed neighbour queue states at each node (a simple estimate of the true empirical distribution), i.e. by changing lines 8-13 in Algorithm 1 to instead sample decision rules for each agent according to the local empirical state distribution, which is verified to be successful. Thus, our approach leans into the centralized training decentralized execution (CTDE) scheme Zhang et al. (2021) and learns a centralized policy, which can then be executed in a decentralized manner among all schedulers. As desired, our approach is applicable to localized queueing systems.
1.4.1 Training on a finite queueing system
We therefore consider alternatives: For one, we could apply a sequential Monte Carlo approach to the problem by simulating $M$ instances of a system from times 00 to some terminal time $T$ that consists of all nodes less than $T$ away from the root node. However, this means that we would have to simulate many finite systems in parallel. Instead, using the fact that the empirical distribution of agent states $\mu^{G_{n},\mathbf{z}_{n}}(t)$ in the finite system converges to $\mathcal{L}(X^{G,x}_{\varnothing}(t))$ as seen in Theorem 1, it should be sufficient to evaluate $\mathcal{L}(X^{G,x}_{\varnothing}(t))$ via the empirical distribution of a sufficiently large system.
1.4.2 Implementation
In order to solve the POMDP, we apply the established proximal policy optimization (PPO) RL method Schulman et al. (2017); Yu et al. (2022) with and without recurrent policies, as commonly and successfully used in POMDP problems Ni et al. (2022). PPO is a policy gradient method with a clipping term in the loss function, such that the policy does not take gradient steps that are too large while learning Schulman et al. (2017); Yu et al. (2022). For our experiments, we have worked with the stable and easy-to-use RLlib 1.10 implementation Liang et al. (2018) of PPO. The overall training code is given in Algorithm 1, which also shows how to analogously apply a trained MFC policy to a finite system. We use diagonal Gaussian neural network policies with $\tanh$-activations and two $256$-node hidden layers, parametrizing MFC MDP actions $\zeta_{t}$ by values in $[0,1]$ for each entry $\zeta_{t}(a\mid z)$, normalized by dividing by the sums $\sum_{a}\zeta_{t}(a\mid z)$.
2 Experiment Setup
In this section, we give an explanation of the different types of graph topologies we have used to verify our aforementioned theoretical analysis. We also give a description of the different load balancing algorithms we have chosen to compare against, as well as discuss hyperparameters used for training and evaluation.
2.1 Topologies
A brief description of the practical topologies of interest, most of which fulfill the convergence in the local weak sense defined earlier, is given here. For each of the following mentioned topologies, the agents are numbered from $1$ to $N$.
First, we consider the simple $1$-D cyclic (CYC-1D) graph, which has extensively been used in the study of queueing networks and is highly local Shortle et al. (2018); Koenigsberg (1982). Each agent $i$ has access to $d=2$ other queues/servers, $\{i-1,i+1\}$, while the edge nodes, $1$ and $N$, form a connection.
Next, we define the cube-connected cycle (CCC) graph. This undirected cubic graph has been considered as a network topology for parallel computing architectures Preparata and Vuillemin (1981); Habibian and Patooghy (2017). It is characterized by the cycle order, $o_{c}$ which is the degree $d$ of each node and defines the total number of nodes $N=o_{c}2^{o_{c}}$ in the graph.
We also apply the torus (TORUS) grid graph that has been repeatedly used to represent distributed systems for parallel computing Deveci et al. (2019); Glantz et al. (2015), as a higher-dimensional extension of the CYC-1D graph. We here consider a $2$-D torus which is a rectangular lattice having $o_{t}$ rows and columns. Every node connects to its $d=4$ nearest neighbours and the corresponding edge nodes are also connected. The total nodes in a $2$-D torus are $N=o_{t}^{2}$.
Another highly general topology is the configuration model (CM). This sophisticated generalized random graph is one of the most important theoretical models in the study of large networks Newman (2018); Fosdick et al. (2018). In contrast to the prior highly local topologies, the CM can capture realistic degree distributions under little clustering. Here every agent is assigned a certain degree, making the graph heterogeneous as compared to every agent having the same degree as in previous mentioned graphs. The degree sequence we have used is in the set $\mathcal{D}=\{2,3\}$ with equal likelihood.
Table 1: System parameters used in the experiments.
@@ 2.1 Topologies
$B$ Queue buffer size $5$
@@
And lastly, for an ablation study on graphical convergence assumptions, we use the Bethe (BETHE) lattice. This cycle-free regular tree graph is used for analysis of many statistical physics, mathematics related models and potential games Ostilli (2012); Szabó and Borsos (2016). It is characterized by a pre-defined lattice depth $o_{b}$, with all nodes in the lattice having the same fixed number of neighbours, $d$. The number of nodes at a depth $o$ away from the root node are given by: $d(d-1)^{(o_{b}-1)}$ and the total nodes in the Bethe lattice are calculated as: $N=1+\sum_{o=1}^{o_{b}}d(d-1)^{(o-1)}$. We use $d=3$ for our experiments. Note that it has already been formally shown in (Lacker et al., 2023, Sections 3.6 and 7.3) that for a sequence of increasing regular trees, the empirical measure does not converge to the same limit law as the root particle, even in the weak sense. This is because even as $N\to\infty$, a large proportion of the nodes are leaves, which greatly influences the behaviour of the empirical measure, as particles at different heights behave differently, and the root particle is the only particle at height zero. We have also verified this mismatch using our experiments in Section 3, though we nevertheless obtain improved performance in certain regimes.
Table 2: Hyperparameter configuration for PPO.
@@
2.2 Load Balancing Algorithms
  • 1.
  • 2.
    MARL-PS (NA-PS): The policy is trained, using PPO with parameter sharing Christianos et al. (2020), as it often gives state-of-the-art performance in various RL benchmarks Yu et al. (2022); Papoudakis et al. (2021), on a smaller number of agents. The learned policy can be used for any arbitrary number of agents. It generates a continuous policy which gives the probability of either sending to your queue or randomly to one of the neighbours queues while only observing the state of your own queue, similar to MF-R.
  • 3.
  • 4.
  • 5.
  • 6.
    Scalable-actor-critic (SAC) Lin et al. (2021): This method learns a discrete policy of sending to any one of the $d$ accessible queues. Each agent learns an individual policy while using as observation the agent’s own queue state and also the queue states of all its neighbours. However, a trained policy cannot be used for an arbitrary number of agents since the policy of an agent is influenced by its neighbours states as well, which is not assumed to be the same for every agent.
Figure 3: (a) JSQ converges as $\Delta t$ increases, validating our simulator. (b) Performance evaluation on a fixed environment while changing only the neural network parameters.
2.3 Training
For all our experiments, we consider that each agent is associated with only one queue, so $N=M$. The servers work at an exponential rate of $\alpha$. At every decision epoch we simulate a Markov modulated arrival rate $\lambda_{t}$, with $\lambda_{0}\sim\mathrm{Unif}(\{\lambda_{h},\lambda_{l}\})$ with transition law for switching between rates
$\mathbb{P}(\lambda_{t+1}=\lambda_{l}\mid\lambda_{t}=\lambda_{h})=0.2,$
$\mathbb{P}(\lambda_{t+1}=\lambda_{h}\mid\lambda_{t}=\lambda_{l})=0.5.$ (2)
Note that these values were chosen to depict the switching of the system between high and low traffic regimes. In principle, any reasonable values can be considered. The rest of the system parameters are given in Table 1. We use the well-established policy gradient based RL algorithm, proximal policy optimization (PPO) as discussed in Section 1.4 with hyperparameters in Table 2. We have used a localized reward function which penalizes for the drops in each queue.
Figure 4: (a): SAC training time to convergence increases with the number of agents, making it difficult to train and use for larger setups. (b): SAC performs worse than our proposed MF-R policy for a number of agents between $5$ and $101$.
Lastly, the number of agents $N$ and degrees $d$ in different graph topologies are fixed during training of the MF-R policies as:
  • 1.
    $1$-D cyclic (CYC-1D): $N=M=101$, $d=2$,
  • 2.
    Cube-connected cycle (CCC): $o_{c}=5$, $N=125$, $d=3$,
  • 3.
    $2$-D Torus grid (TORUS): $o_{t}=11$, $N=111$, $d=4$,
  • 4.
    Configuration model (CM): $N=101$, $d\in\{2,3\}$,
  • 5.
    Bethe lattice (BETHE): $o_{b}=5$, $N=94$, $d=3$.
Figure 5: Performance comparison of the learned MF-R policy to NA-PS, JSQ, RND and OWN algorithms, on a $1$-D cyclic graph, over a range of synchronization delays is shown, with $95\%$ confidence intervals depicted by error bars. The degree of each agent is $d=2$ and the number of agents (queues) used to make the graph are $N\in\{9,21,91,901,3501,5001\}$. It can be seen that for the in-between range of $\Delta t$ our MF-R policy is the optimal one.
3 Experiment Results
We now present the performance comparison of the load balancing policies of Section 2.2 on graph topologies from Section 2.1. For the exact simulation of associated continuous-time Markov chains $y$, we sample exponential waiting times of all events using the Gillespie algorithm Gillespie (1977). For training, we use the same simulation horizon for each episode consisting of $T=50$ discrete decision epochs, and for comparability the performance is evaluated in terms of the average number of packets dropped per $50$ time units and queue. However, note that simulating the same time span with different $\Delta t$ does provide slightly different results, due to the switching between high and low traffic regimes after each $\Delta t$ epoch. Each evaluation was repeated for $100$ simulated episodes, and error bars in all figures depict the $95\%$ confidence interval.
Figure 6: Performance of the MF-R policy for increasingly large CYC-1D graphs. The red horizontal line indicates the evaluated episode return of the learned MF-R policy during training on $N=101$, (MF-MFC). Shaded regions depict the $95\%$ confidence intervals. As the system size increases, the performance of the learned policy (black) increasingly converges to the mean field system performance (red), validating the accuracy of our mean field formulation and choice of $N$ for training.
Figure 7: Performance over $\Delta t\in\{1,2,\ldots,10\}$ for large sparse graphs with underlying topologies of CYC-1D in (a), CCC in (b), TORUS in (c) and CM in (d). The number of nodes used to generate the graphs is given at the top of each subfigure. It can be seen that for smaller synchronization delays, JSQ is the optimal policy, while for higher delays RND is the optimal one as expected from prior literature Mitzenmacher (2001). For the in-between range of delays, our MF-R policy performs the best and approaches the optimal performance of RND as the $\Delta t$ increases. OWN and RND have equivalent performance in CYC, CCC and TORUS, since fixed degrees imply the same packet arrival rate when always sending to your own queue, or always sending randomly to any neighbour.
Figure 8: Performance comparison on large-sized Bethe lattice graph. The MF-R can be worse than RND due to violation of modelling assumptions.
Simulator analysis

### 22 Mar 2024 v2

@@ 2.4. Reinforcement Learning
In our experiments we also allow to simply insert the empirical distribution of the locally observed neighbour queue states at each node (a simple estimate of the true empirical distribution), to instead sample decision rules for each agent according to the local empirical state distribution, which is verified to be successful. Thus, our approach leans into the centralized training decentralized execution scheme (Zhang et al., 2021) and learns a centralized policy, which can then be executed in a decentralized manner among all schedulers. As desired, our approach is applicable to localized queueing systems.
In order to solve the POMDP, we apply the established proximal policy optimization (PPO) RL method (Schulman et al., 2017; Yu et al., 2022) with and without recurrent policies, as commonly and successfully used in POMDP problems (Ni et al., 2022). PPO is a policy gradient method with a clipping term in the loss function, such that the policy does not take gradient steps that are too large while learning (Schulman et al., 2017; Yu et al., 2022). The overall training algorithm is given in Appendix, which also shows how to analogously apply a trained MFC policy to a finite system.
2.4.1. Training on a finite queueing system
We therefore consider alternatives: For one, we could apply a sequential Monte Carlo approach to the problem by simulating $M$ instances of a system from times 00 to some terminal time $T$ that consists of all nodes less than $T$ away from the root node. However, this means that we would have to simulate many finite systems in parallel. Instead, using the fact that the empirical distribution of agent states $\mu^{G_{n},\mathbf{z}_{n}}(t)$ in the finite system converges to $\mathcal{L}(X^{G,x}_{\varnothing}(t))$ as seen in Theorem 2.1, it should be sufficient to evaluate $\mathcal{L}(X^{G,x}_{\varnothing}(t))$ via the empirical distribution of a sufficiently large system.
3. Experiment Setup
In this section, we give an explanation of the different types of graph topologies we have used to verify our aforementioned theoretical analysis. We also give a description of the different used load balancing algorithms.
3.1. Topologies
A brief description of the practical topologies of interest, most of which fulfill the convergence in the local weak sense defined earlier is given here, where the agents are numbered from $1$ to $N$. (i) First, we consider the simple CYC-1D graph, which has extensively been used in the study of queueing networks and is highly local (Shortle et al., 2018). Each agent $i$ has access to $d=2$ other queues/servers, $\{i-1,i+1\}$, while the edge nodes, $1$ and $N$, form a connection. (ii) Next, we define the cube-connected cycle (CCC) graph. This undirected cubic graph has been considered as a network topology for parallel computing architectures (Habibian and Patooghy, 2017). It is characterized by the cycle order, $o_{c}$ which is the degree $d$ of each node and defines the total number of nodes $N=o_{c}2^{o_{c}}$ in the graph. (iii) We also apply the torus (TORUS) grid graph that has been repeatedly used to represent distributed systems for parallel computing (Deveci et al., 2019), as a higher-dimensional extension of the CYC-1D graph. We here consider a $2$-D torus which is a rectangular lattice having $o_{t}$ rows and columns. Every node connects to its $d=4$ nearest neighbours and the corresponding edge nodes are also connected. The total nodes in a $2$-D torus are $N=o_{t}^{2}$. (iv) Another highly general topology is the configuration model (CM). This sophisticated generalized random graph is one of the most important theoretical models in the study of large networks (Fosdick et al., 2018). In contrast to the prior highly local topologies, the CM can capture realistic degree distributions under little clustering. Here every agent is assigned a certain degree, making the graph heterogeneous as compared to every agent having the same degree as in previous mentioned graphs. The degree sequence we have used is in the set $\mathcal{D}=\{2,3\}$ with equal likelihood. (v) And lastly, for an ablation study on graphical convergence assumptions, we use the Bethe (BETHE) lattice. This cycle-free regular tree graph is used for analysis of many statistical physics, mathematics related models and potential games (Szabó and Borsos, 2016). It is characterized by a pre-defined lattice depth $o_{b}$, with all nodes in the lattice having the same fixed number of neighbours, $d$. The number of nodes at a depth $o$ away from the root node are given by: $d(d-1)^{(o_{b}-1)}$ and the total nodes in the Bethe lattice are calculated as: $N=1+\sum_{o=1}^{o_{b}}d(d-1)^{(o-1)}$. We use $d=3$ for our experiments. Note that it has already been formally shown in (Lacker et al., 2023, Sections 3.6 and 7.3) that for a sequence of increasing regular trees, the empirical measure does not converge to the same limit law as the root particle, even in the weak sense. This is because even as $N\to\infty$, a large proportion of the nodes are leaves, which greatly influences the behaviour of the empirical measure, as particles at different heights behave differently, and the root particle is the only particle at height zero. We have also verified this mismatch using our experiments in Section 4, though we nevertheless obtain improved performance in certain regimes.
3.2. Load Balancing Algorithms
  • MARL-PS (NA-PS): The policy is trained, using PPO with parameter sharing (Christianos et al., 2020), as it often gives state-of-the-art performance in various RL benchmarks (Yu et al., 2022; Papoudakis et al., 2021), on a smaller number of agents. The learned policy can be used for any arbitrary number of agents. It generates a continuous policy which gives the probability of either sending to your queue or randomly to one of the neighbours queues while only observing the state of your own queue, similar to MF-R.
Figure 2. Performance comparison of the learned MF-R policy to NA-PS, JSQ, RND and OWN algorithms, on a CYC-1D graph, over a range of $\Delta t$s is shown, with $95\%$ confidence intervals depicted by error bars. The degree of each agent is $d=2$ and the number of agents (queues) used to make the graph are $N\in\{9,21,91,901,3501,5001\}$.
3.3. Training
For all our experiments, we consider that each agent is associated with only one queue, so $N=M$. The servers work at an exponential rate of $\alpha$. At every decision epoch we simulate a Markov modulated arrival rate $\lambda_{t}$, with $\lambda_{0}\sim\mathrm{Unif}(\{\lambda_{h},\lambda_{l}\})$ with transition law for switching between rates: $\mathbb{P}(\lambda_{t+1}=\lambda_{l}\mid\lambda_{t}=\lambda_{h})=0.2,$ and $\mathbb{P}(\lambda_{t+1}=\lambda_{h}\mid\lambda_{t}=\lambda_{l})=0.5$. Note that these values were chosen to depict the switching of the system between high and low traffic regimes. In principle, any reasonable values can be considered. We use the well-established policy gradient based RL algorithm, proximal policy optimization (PPO) with a localized reward function which penalizes for the drops in each queue. The rest of the system parameters and the PPO hyperparameters are given in Appendix. Lastly, the number of agents $N$ and degrees $d$ in different graph topologies are fixed during training of the MF-R policies as:
  • $1$-D cyclic (CYC-1D): $N=M=101$, $d=2$,
  • Cube-connected cycle (CCC): $o_{c}=5$, $N=125$, $d=3$,
  • $2$-D Torus grid (TORUS): $o_{t}=11$, $N=111$, $d=4$,
  • Configuration model (CM): $N=101$, $d\in\{2,3\}$,
  • Bethe lattice (BETHE): $o_{b}=5$, $N=94$, $d=3$.
4. Experiment Results
We now present the performance comparison of the load balancing policies of Section 3.2 on graph topologies from Section 3.1. For the exact simulation of associated continuous-time Markov chains $y$, we sample exponential waiting times of all events using the Gillespie algorithm. For training, we use the same simulation horizon for each episode consisting of $T=50$ discrete decision epochs, and for comparability the performance is evaluated in terms of the average packet drops which is calculated as the sum over $T$ decision epochs of the average number of total packets dropped in all queues. However, note that simulating the same time span with different $\Delta t$ does provide slightly different results, due to the switching between high and low traffic regimes after each $\Delta t$ epoch. Each evaluation was repeated for $100$ simulated episodes, and error bars depict the $95\%$ confidence interval.
Finite system performance
The second result of interest is the performance analysis on the CYC-1D graph over a range of $\Delta t$ for the learned MF-R policy as compared to the NA-PS, JSQ, RND and OWN algorithms, see Figure 2. The number of agents (queues) used to generate the graph are $N\in\{9,21,91,901,3501,5001\}$, with every agent always having the fixed degree $d=2$. For this graph, we also trained an NA-PS policy on a sufficiently small size with $N=M=5$ to obtain convergence even with batch and minibatch sizes of $50000$ and $8000$. We used PPO with parameter sharing so that the learned policy could be evaluated on any graph size.
It can be seen that JSQ is the best strategy at $\Delta t=1$, while the performance of MF-R is very close to JSQ. And as expected, the performance of JSQ deteriorates with increasing synchronisation delay $\Delta t$. For an intermediate value of $\Delta t\in\{3,5,7\}$, neither JSQ nor RND are the optimal load balancing policies. Here, our learned MF-R policy performs best and approaches the optimal RND performance as $\Delta t$ increases to $10$. As discussed earlier, it has already been proven in certain scenarios that as $\Delta t\to\infty$, RND is the optimal load balancing policy (Mitzenmacher, 2001). The learned NA-PS policy has a similar trend to our MF-R policy, but is unable to outperform it even for small systems. For the CYC-1D, CCC, and TORUS graphs with homogeneous degrees, the performance of OWN in the limit – sending only to its own queue – is equivalent to RND, since for homogeneous degrees, sending to your own queue results in the same packet arrival rate at each queue as randomly sending to any queue in your neighbourhood. In Figure 3, it can also be seen that as the size of the graph increases, the performance of the learned policy (black) increasingly converges to the mean field system performance (red), validating the accuracy of our mean field formulation and choice of $N$ for training.
Performance in large systems
In Figure 4 each subfigure shows results for a different topology. We show the performance over a range of $\Delta t\in\{1,2,\ldots,10\}$ for large sparse networks. Figure 4(a) is for the CCC graph with $d=3$ and $N=4608$. Figure 4(b) shows the performance for the TORUS graph with $d=4$, $o_{t}=70$ and $N=4900$, while Figure 4(c) is for the CM graph for $N=5001$ with uniformly distributed degrees in the set $d\in\{2,3\}$. Our learned MF-R policy is compared to JSQ, RND and OWN algorithms. The following observations can be made for all the subfigures: (i) For small $\Delta t=1$ JSQ performs better than all the algorithms which is theoretically guaranteed, however performance of MF-R is very close to it; (ii) For in between range of $\Delta t=\{2,3,\ldots,7\}$ MF-R has the best performance in terms of packets dropped; (iii) Except in the CM graph (where the varying degrees lead to differing packet arrival rates between OWN and RND), we find that OWN and RND again have equivalent performance by regularity of the graph, as discussed in the previous paragraph; (iv) For higher $\Delta t\in\{8,9,10\}$ our MF-R policy performance coincides with the RND policy, where the RND policy has been theoretically proven to be optimal as the synchronization delay $\Delta t\to\infty$ (Mitzenmacher, 2001). Hence, we find that MF-R policy performs consistently well in various topologies.
Figure 3. Performance of the MF-R policy for increasingly large CYC-1D graphs. The red horizontal line indicates the evaluated episode return of the learned MF-R policy during training on $N=101$, (MF-MFC). Shaded regions depict the $95\%$ confidence intervals.
Figure 4. Performance over $\Delta t\in\{1,2,\ldots,10\}$ for large sparse graphs with underlying topologies of CCC in (a), TORUS in (b) and CM in (c). The number of nodes used to generate the graphs is given at the top of each subfigure.
Ablation on topological assumptions
As mentioned in Section3.1, the theoretical analysis and work in (Lacker et al., 2023) does not apply to the Bethe topology, and the effect is further supported by our experimental results. Figure 5 shows the performance comparison for Bethe lattice graphs of the learned MF-R policy with JSQ, RND, and OWN algorithms over different $\Delta t$. We show the result only for a large Bethe network, with $o_{b}=11$, $d=3$ and $N=6142$. OWN outperforms RND more significantly than in CM, since disproportionately many packets arrive at nodes connected to leaf nodes. It can be seen that the system behaves as expected when $\Delta t$ is small, with JSQ quickly outperformed by MF-R, while MF-R eventually performs worse than OWN as $\Delta t$ increases. This behaviour is due to the fact that in structures such as regular trees, the central node as the root node is not sufficient to represent the performance of all nodes, especially the many leaf nodes that behave differently even in the limit. Because of this loss of regularity, theoretical guarantees fail when scaling large Bethe or similar topologies, as explained in Section 3.1. Even so, we see that the learned policy can outperform existing solutions in certain regimes, e.g., at $\Delta t=3$.
For the sake of completeness we have also provided the results of following experiments in Appendix: (i) convergence analysis of our simulator, (ii) effect of different neural network parameters, (iii) two other MF based policies (iv) comparison to the SAC algorithm (Li, 2021), (v) effect of increasing buffer size, (vi) having heterogenous servers, (vi) using recurrent neural networks with PPO and (vii) use of different observation spaces for MF policies.
Figure 5. Performance comparison on large-sized Bethe lattice graph. The MF-R can be worse than RND due to violation of modelling assumptions.
5. Conclusion
Overall, we have learned efficient algorithms for load balancing in highly local queuing systems with some degree of theoretical guarantees through sparse mean-field approximations. The approach has been positively evaluated in comparison to well-known baselines, and can scale to queueing systems of arbitrary size with arbitrary wireless communication delays. Future work could attempt to quantify rates of convergence for approximate optimality guarantees of MFC policies under the framework of (Lacker et al., 2023), or consider dynamic programming principles. One could also look into the analysis of graph structures such as Bethe and regular trees, for which the current sparse MFC modelling does not suffice.
Appendix A Appendix
This Appendix contains the proofs, algorithm, hyperparameter values and additional experiments connected to the paper titled: ”Sparse Mean Field Load Balancing in Large Localized Queueing Systems”.
Appendix B Proof
Proof of Theorem 2.1.
We apply the framework of (Lacker et al., 2023), which does not include actions or rewards, by including actions and rewards via separate time steps. Specifically, each time step is split in three, and we define the agent state space $\mathcal{X}\coloneqq\mathcal{Z}\cup(\mathcal{Z}\times\mathcal{A})\cup(\mathcal{Z}\times\mathcal{A}\times[0,D_{\mathrm{max}}])$ for each of the agents, indicating at any third decision epoch $t=0,3,6,\ldots$ the state of its queue. After each such epoch, the agent states will contain the state of its queue together with its choice of action at times $t=1,4,\ldots$, and lastly also the number of packet drops at times $t=2,5,\ldots$. Here, $D_{\mathrm{max}}$ is the maximum expected number of packet drops and is given by the $(d+1)$ times the maximum per-scheduler arrival rate $\lambda_{\mathrm{max}}=\lambda_{h}$.
We formally rewrite the system defined earlier, using the symbol $\mathbf{x}$ for the state of the rewritten system instead of $\mathbf{z}$. As a result, each agent $i$ is endowed with a random local $\mathcal{X}$-valued state variable $X^{G,x}_{i}(t)$ at each time $t$. For the dynamics, we let $S^{\sqcup}(\mathcal{X})$ denote the space of unordered terminating sequences (up to some maximum degree) with the discrete topology as in (Lacker et al., 2023), and define in the following a system dynamics function $F\colon\mathcal{X}\times S^{\sqcup}(\mathcal{X})\times\Xi\to\mathcal{X}$ returning a new state for any current local state $X^{G,x}_{i}(t)$, any current $S^{\sqcup}(\mathcal{X})$-valued neighbourhood queue fillings $X^{G,x}_{N_{i}}(t)$ and i.i.d. sampled $\Xi$-valued noise $\xi_{i}(t+1)$. More precisely, we use
$X^{G,x}_{i}(t+1)=F[X^{G,x}_{i}(t),X^{G,x}_{N_{i}}(t),\xi_{i}(t+1)]$
For the first step, fixing any $\pi$, we let
$\xi_{i,z}\sim\mathrm{Bernoulli}(\pi_{\lfloor t/3\rfloor}(z)),$
for all $z\in\mathcal{Z}$, to sample action from $\pi$. Deferring for a moment the second step, for the new state in the third step, we use the Kolmogorov forward equation for the queue state during the epoch for any initial $(x,\mu)$:
$P(x,\mu)=\exp{\mathbf{Q}(x,\mu)}\cdot e_{x_{1}}\in\Delta^{\mathcal{Z}}$
with unit vector $e_{x_{1}}$, which results from Poisson thinning, as the equivalent arrival rate at a queue of an agent currently in state $x$ with neighbourhood $\mu$ will be given by $\lambda\left(1-x_{2}+\frac{\mu(\mathbf{1}_{\mathcal{Z}\times\{1\}\times S^{\sqcup}(\mathcal{X})})}{d}\right)$. Thus, we have the transition rate matrix $\mathbf{Q}(x,\mu)$ with $\mathbf{Q}_{i,i-1}=\lambda\left(1-x_{2}+\frac{\mu(\mathbf{1}_{\mathcal{Z}\times\{1\}\times S^{\sqcup}(\mathcal{X})})}{d}\right)$, $\mathbf{Q}_{i-1,i}=\alpha$ for $i=1,\ldots,B$, and where $\mathbf{Q}_{i,i}=-\sum_{j}\mathbf{Q}(\lambda,z)_{j,i}$ for $i=0,\ldots,B$, and zero otherwise. Therefore, the next queue filling is sampled as
$\xi_{i,x,\mu}\sim\mathrm{Categorical}(P(x,\mu)).$
Lastly, for the second step we consider the non-random conditional expectation of packet losses during any epoch:
$\xi_{i,x_{1},x_{2},\mu}=[\exp{\bar{\mathbf{Q}}(x,\mu)}\cdot e_{x_{1}}]_{B+1}$
under the executed actions (allocations), analogously to the new state, by adding another row to the transition matrix $\mathbf{Q}(x,\mu)$, giving $\bar{\mathbf{Q}}(x,\mu)$ where $\bar{\mathbf{Q}}_{B+1,B}=\lambda\left(1-x_{2}+\frac{\mu(\mathbf{1}_{\mathcal{Z}\times\{1\}})}{d}\right)$, i.e. counting all the expected packet arrivals whenever the queue is already full ($B$).
(1) $\mu^{G_{n},\mathbf{z}_{n}}(t)\to\mathcal{L}(X^{G,x}_{\varnothing}(t))\quad\text{in distribution in $\mathcal{P}(\mathcal{X})$}$
since we split any time step into three by the prequel, the continuous mapping theorem, and dominated convergence, where we use the continuous function
$r^{\infty}(\mu)=\int\mathbf{x}_{3}1_{x\in\mathcal{Z}\times\mathcal{A}\times[0,D_{\mathrm{max}}]}\mu(\mathrm{d}x),$
where $\mu=\mathcal{P}(\mathcal{X})$, to sum up the expected packet drops in the system, since the integrand is continuous and bounded (under the sum topology for the union in $\mathcal{X}$, and the product topology for products in $\mathcal{X}$). ∎
Proof of Corollary 2.2.
Define the non-zero (by assumption) optimality gap:
$\Delta J\coloneqq J^{G,\mathbf{z}}(\pi_{i})-\max_{j\neq i}J^{G,\mathbf{z}}(\pi_{j})>0.$
Then, by Theorem 2.1, there exists $n$ such that:
$\max_{i}\left|J^{G_{n},\mathbf{z}_{n}}(\pi_{i})-J^{G,\mathbf{z}}(\pi_{i})\right|\leq\frac{\Delta J}{2}.$
Therefore,
is the desired conclusion. ∎
Appendix C Algorithm
In order to solve the POMDP, we apply the established proximal policy optimization (PPO) RL method (Schulman et al., 2017; Yu et al., 2022) with and without recurrent policies, as commonly and successfully used in POMDP problems (Ni et al., 2022). PPO is a policy gradient method with a clipping term in the loss function, such that the policy does not take gradient steps that are too large while learning (Schulman et al., 2017; Yu et al., 2022). For our experiments, we have worked with the stable and easy-to-use RLlib 1.10 implementation (Liang et al., 2018) of PPO. The overall training code is given in Algorithm 1 given in , which also shows how to analogously apply a trained MFC policy to a finite system. We use diagonal Gaussian neural network policies with $\tanh$-activations and two hidden layers, parametrizing MFC MDP actions $\zeta_{t}$ by values in $[0,1]$ for each entry $\zeta_{t}(a\mid z)$, normalized by dividing by the sums $\sum_{a}\zeta_{t}(a\mid z)$.
@@ Appendix C Algorithm
The rest of the system parameters and the PPO hyperparameter are given here in Tables 1 and 2, respectively.
Table 1. System parameters used in the experiments.
@@
$B$ Queue buffer size $5-20$
@@
Table 2. Hyperparameter configuration for PPO.
@@
Appendix D Additional Experiments
Figure 6. (a) JSQ converges as $\Delta t$ increases, validating our simulator. (b) Performance evaluation on a fixed environment while changing only the neural network parameters.
D.1. Simulator analysis